Show complete solution. Use equation function.

1)Find the general solution to the equation (cos y + 2) dy/dx = 2x.


1
Expert's answer
2021-06-02T10:30:24-0400
"(\cos y+2)dy/dx=2x"

"(\cos y+2)dy=2xdx"

"\int(\cos y+2)dy=\int2xdx"

"\sin y+2y=x^2+C"


"\sin y+2y-x^2=C"